# On Marginal Stability in Low Temperature Spherical Spin Glasses

### v2

@@ 3 Alternate Proof of Marginal Stability for Generic Models
(Note the similarity to the “even generic” hypothesis of Corollary 1.7, which would also suffice here.) We prove the following, which is a special case of (1.13) in Corollary 1.4.
@@
Suppose $\xi$ is generic and not quadratic, and $1$ is not an isolated point in $T$ (recall (1.25)). Then for any $\varepsilon>0$, if ${\boldsymbol{\sigma}}_{\beta}\sim\mu_{\beta}$ for $\beta\geq\beta_{*}(\xi,\varepsilon)$ sufficiently large, and $\delta,c$ are small depending on $(\xi,\varepsilon,\beta)$, and $N$ is sufficiently large, we have with probability $1-e^{-cN}$:
@@
Let ${\cal M}$ denote the space of all right-continuous non-decreasing functions $x:[0,1]\to[0,1]$ with $x(\hat{q})=1$ for some $\hat{q}<1$ (which may depend on $x$). Let $\hat{x}(q)=\int_{q}^{1}x(s)\leavevmode\nobreak\ {\rm d}s$ and define the Crisanti–Sommers functional
${\mathcal{P}}(x;\xi)=\frac{1}{2}\left\{\xi^{\prime}(0)\hat{x}(0)+\int_{0}^{1}\xi^{\prime\prime}(q)\hat{x}(q)\leavevmode\nobreak\ {\rm d}q+\int_{0}^{\widehat{q}}\frac{{\rm d}q}{\hat{x}(q)}+\log(1-{\widehat{q}})\right\}.$ (3.3)
Note that $\hat{x}(q)=1-q$ for $q>{\widehat{q}}$, so this functional is independent of ${\widehat{q}}$. The spherical Parisi formula at positive temperature is as follows.
@@
$F(\beta)$ $\equiv\operatorname*{p-lim}_{N\to\infty}F_{N,\beta}=\inf_{(\zeta,b)}{\mathcal{P}}_{\beta}(\zeta,b),$
${\mathrm{GS}}$ $=\lim_{\beta\to\infty}F(\beta)/\beta.$
Moreover the minimizers $(\zeta_{\beta},b_{\beta})$ exist and are unique. With $(\zeta,L)$ the zero-temperature minimizers, one has
@@
There is a positive-temperature analog of Proposition 1.13 characterizing $(\zeta_{\beta},b_{\beta})$, but this will not be needed. We instead use the following qualitative result on the overlaps between Gibbs samples. Here and below, ${\rm supp}(\zeta_{\beta})\subseteq[0,1)$ is the set of points of increase of $\zeta_{\beta}$.
@@
As stated in (3.2), the zero-temperature optimizer $\zeta$ is the vague limit as $\beta\to\infty$ of $\beta\zeta_{\beta}$. Hence it suffices to show that $\zeta$ is strictly increasing in a neighborhood of $1$. Proposition 1.13 implies that $T$ contains a non-trivial interval $(1-\varepsilon_{*}(\xi),1)$. On this interval, we must have $g^{\prime\prime}(q)=0$, which easily rearranges to
The latter is a analytic function on an open complex neighborhood of $[1-\varepsilon_{*},1]$ and is not identically zero since $\xi$ is not quadratic. Thus it has finitely many zeros; this completes the proof. ∎
We consider the event that a large number $K$ of Gibbs samples have all overlaps approximately equal to $q$:
@@ 3.1 Preparatory Lemmas
$\langle x_{i}-x_{j},x_{i}-x_{k}\rangle=\frac{\|x_{i}-x_{j}\|^{2}+\|x_{i}-x_{k}\|^{2}-\|x_{j}-x_{k}\|^{2}}{2}=\frac{1}{2}\pm O(\varepsilon).$
@@
$\langle y_{i},y_{i}\rangle=(d+2)^{-2}\sum_{k,\ell=1}^{d+2}\langle x_{i}-x_{k},x_{i}-x_{\ell}\rangle\\ =\left(\frac{d+1}{d+2}\right)^{2}\pm O(\varepsilon).$
@@
$\langle y_{i},y_{j}\rangle=(d+2)^{-2}\sum_{k,\ell=1}^{d+2}\langle x_{i}-x_{k},x_{j}-x_{\ell}\rangle=-\frac{d+1}{(d+2)^{2}}\pm O(\varepsilon).$
@@
${\boldsymbol{v}}={\boldsymbol{v}}_{\parallel}+{\boldsymbol{v}}_{\perp}+R({\boldsymbol{\sigma}},v){\boldsymbol{\sigma}}$ (3.6)
with $({\boldsymbol{v}}_{\parallel},{\boldsymbol{v}}_{\perp})\in U_{K}\times U_{K}^{\perp}$. Note that if ${\boldsymbol{v}}={\boldsymbol{\sigma}}^{\prime}-{\boldsymbol{\sigma}}$ then ${\boldsymbol{v}}_{\parallel}+{\boldsymbol{v}}_{\perp}$ is proportional to the derivative $\gamma^{\prime}(0)$ for $\gamma$ a geodesic path from ${\boldsymbol{\sigma}}$ to ${\boldsymbol{\sigma}}^{\prime}$. We also set $U_{K,\lambda}$ to be the $\lambda\sqrt{N}$-neighborhood of $U_{K}({\boldsymbol{\sigma}})+{\boldsymbol{\sigma}}$ in ${\mathbb{R}}^{N}$ (note that $U_{K}\subseteq{\mathcal{T}}({\boldsymbol{\sigma}})\subseteq{\mathbb{R}}^{N}$ can be naturally viewed as a subset of ${\mathbb{R}}^{N}$).
@@
This amounts to a second order Taylor expansion of $H_{N}$ along the geodesic path from ${\boldsymbol{\sigma}}$ to ${\boldsymbol{\sigma}}^{\prime}$. The required estimate on the third derivative holds since $H_{N}\in K_{N}$. ∎
@@
$\leq H_{N}({\boldsymbol{\sigma}})+O(\eta\varepsilon^{1/2}+\varepsilon^{3/2})N-C\|{\boldsymbol{v}}^{\perp}\|^{2}/2\leq H_{N}({\boldsymbol{\sigma}})-C\lambda^{2}N/3.$
@@
$\|{\boldsymbol{v}}^{i}_{\parallel}-{\boldsymbol{v}}^{j}_{\parallel}\|$ $=\|{\boldsymbol{v}}^{i}-{\boldsymbol{v}}^{j}\|\pm O(\|{\boldsymbol{v}}^{i}-{\boldsymbol{v}}^{i}_{\parallel}\|+\|{\boldsymbol{v}}^{j}-{\boldsymbol{v}}^{j}_{\parallel}\|)=\|{\boldsymbol{\sigma}}^{i}-{\boldsymbol{\sigma}}^{j}\|\pm O(\|{\boldsymbol{v}}^{i}-{\boldsymbol{v}}^{i}_{\parallel}\|+\|{\boldsymbol{v}}^{j}-{\boldsymbol{v}}^{j}_{\parallel}\|)$
$=\Big{(}\sqrt{2-2(1-\varepsilon)^{2}}\pm O(\delta\varepsilon^{-1/2}+\lambda)\Big{)}\sqrt{N}.$
@@
The main part of the proof is the lower bound on ${\boldsymbol{\lambda}}_{K}(\nabla^{2}_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}_{\beta}))$. Suppose for sake of contradiction that (3.5) holds for some $K,C>0$, and let $S_{{\rm eigen}}$ denote the event that ${\boldsymbol{\lambda}}_{K}(\nabla^{2}_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}_{\beta}))\leq-C$. Choose small $\varepsilon$ depending on $(K,C)$ such that $1-\varepsilon\in{\rm supp}(\zeta)$. Let $\eta\ll\delta$ be small depending on $(\xi,K,C,\varepsilon,\beta)$, and define for i.i.d. Gibbs samples $({\boldsymbol{\sigma}}_{\beta}^{1},\dots,{\boldsymbol{\sigma}}_{\beta}^{K+3})$ the event
@@
In light of Lemma 3.7 and the assumption $S_{{\rm bounded}}$, we find that $S_{{\rm eigen}}$ indeed cannot hold. Finally it follows from (3.7) that ${\mathbb{P}}[S_{{\rm eigen}}]\leq\frac{\varepsilon}{2}+e^{-cN}\leq\varepsilon$ which concludes the proof. ∎
@@ Acknowledgements
Thanks to Brice Huang for helpful discussions, and for comments on a previous version of the paper. We also thank Andrea Montanari, Pierfrancesco Urbani, and Lenka Zdeborová for enlightening conversations.

### v3

@@ 3 Alternate Proof of Marginal Stability for Generic Models
(Note the similarity to the “even generic” hypothesis of Corollary 1.7, which would also suffice here.) We prove the following, which is a special case of (1.9) and (1.14).444 For instance the hypothesis is slightly stronger than our previous full RSB assumption by Proposition 3.1, and ${\boldsymbol{\sigma}}_{\beta}$ is a near-ground state with high probability for large $\beta$ by Lemma 1.12.
@@
Suppose $\xi$ is generic, and $1$ is not an isolated point in $T$ (recall (1.27)). Then for any $\varepsilon>0$, if ${\boldsymbol{\sigma}}_{\beta}\sim\mu_{\beta}$ for $\beta\geq\beta_{*}(\xi,\varepsilon)$ sufficiently large, and $\delta,c$ are small depending on $(\xi,\varepsilon,\beta)$, and $N$ is sufficiently large, we have with probability at least $1-\varepsilon$:
@@
Let ${\cal M}$ denote the space of all right-continuous non-decreasing functions ${\zeta_{\beta}}:[0,1]\to[0,1]$ with ${\zeta_{\beta}}(\hat{q})=1$ for some $\hat{q}<1$ (which may depend on ${\zeta_{\beta}}$). (Here we have in mind the positive temperature setting $\beta<\infty$, but the definition of the space ${\cal M}$ does not depend on $\beta$.) Let ${\hat{\zeta}_{\beta}}(q)=\int_{q}^{1}{\zeta_{\beta}}(s)\leavevmode\nobreak\ {\rm d}s$ and define the Crisanti–Sommers functional
${\mathcal{P}}({\zeta_{\beta}};\xi)=\frac{1}{2}\Big{(}\xi^{\prime}(0){\hat{\zeta}_{\beta}}(0)+\int_{0}^{1}\xi^{\prime\prime}(q){\hat{\zeta}_{\beta}}(q)\leavevmode\nobreak\ {\rm d}q+\int_{0}^{\widehat{q}}\frac{{\rm d}q}{{\hat{\zeta}_{\beta}}(q)}+\log(1-{\widehat{q}})\Big{)}.$ (3.3)
Note that ${\hat{\zeta}_{\beta}}(q)=1-q$ for $q>{\widehat{q}}$, so this functional is independent of ${\widehat{q}}$. The spherical Parisi formula at positive temperature is as follows.
@@
$F(\beta)$ $\equiv\operatorname*{p-lim}_{N\to\infty}F_{N,\beta}={\inf_{\zeta_{\beta}\in{\cal M}}{\mathcal{P}}(\zeta_{\beta};\beta\xi)},$
${\mathrm{GS}}(\xi)$ $=\lim_{\beta\to\infty}F(\beta)/\beta.$
Further a unique minimizer exists, which we also denote by $\zeta_{\beta}$. With $(\zeta,L)$ the zero-temperature minimizers, one has
@@
There is a positive-temperature analog of Proposition 1.13 characterizing ${\zeta_{\beta}}$, but this will not be needed. We instead use the following qualitative result on the overlaps between Gibbs samples. Recall that ${\rm supp}(\zeta_{\beta})\subseteq[0,1)$ is the set of points of increase of $\zeta_{\beta}$.
@@
As stated in (3.2), the zero-temperature optimizer $\zeta$ is the vague limit as $\beta\to\infty$ of $\beta\zeta_{\beta}$. Hence it suffices to show the first claim that $\zeta$ is strictly increasing in a neighborhood of $1$, which is shown in Corollary 1.14 (since $\xi$ is generic it cannot be quadratic). ∎
Next we define the event that a large constant number $K$ of Gibbs samples have all overlaps approximately $q$:
@@
$\langle x_{i}-x_{j},x_{i}-x_{k}\rangle=\frac{\|x_{i}-x_{j}\|^{2}+\|x_{i}-x_{k}\|^{2}-\|x_{j}-x_{k}\|^{2}}{2}=\frac{1}{2}{+}O(\varepsilon).$
@@
$\langle y_{i},y_{i}\rangle=(d+2)^{-2}\sum_{k,\ell=1}^{d+2}\langle x_{i}-x_{k},x_{i}-x_{\ell}\rangle\\ =\left(\frac{d+1}{d+2}\right)^{2}{+}O(\varepsilon).$
@@
$\langle y_{i},y_{j}\rangle=(d+2)^{-2}\sum_{k,\ell=1}^{d+2}\langle x_{i}-x_{k},x_{j}-x_{\ell}\rangle=-\frac{d+1}{(d+2)^{2}}{+}O(\varepsilon).$
@@
${\boldsymbol{v}}={\boldsymbol{v}}_{\parallel}+{\boldsymbol{v}}_{\perp}+R({\boldsymbol{\sigma}},{\boldsymbol{v}}){\boldsymbol{\sigma}}$ (3.6)
with $({\boldsymbol{v}}_{\parallel},{\boldsymbol{v}}_{\perp})\in U_{K}\times U_{K}^{\perp}$. Note that if ${\boldsymbol{v}}={\boldsymbol{\sigma}}^{\prime}-{\boldsymbol{\sigma}}$ then ${\boldsymbol{v}}_{\parallel}+{\boldsymbol{v}}_{\perp}$ is proportional to the derivative $\gamma^{\prime}(0)$ for $\gamma$ a geodesic path from ${\boldsymbol{\sigma}}$ to ${\boldsymbol{\sigma}}^{\prime}$. We also set $U_{K,\lambda}$ to be the $\lambda\sqrt{N}$-neighborhood of $U_{K}({\boldsymbol{\sigma}})+{\boldsymbol{\sigma}}$ in ${\mathbb{R}}^{N}$ (note that $U_{K}\subseteq{\mathcal{T}}({\boldsymbol{\sigma}})\subseteq{\mathbb{R}}^{N}$ can be naturally viewed as a linear subspace of ${\mathbb{R}}^{N}$).
@@
Recalling (1.24), this amounts to a second order Taylor expansion of $H_{N}$ along the geodesic path from ${\boldsymbol{\sigma}}$ to ${\boldsymbol{\sigma}}^{\prime}$. The required estimate on the third derivative holds since $H_{N}\in K_{N}$. ∎
@@
$\leq H_{N}({\boldsymbol{\sigma}})+O(\eta\varepsilon^{1/2}+\varepsilon^{3/2})N-C\|{\boldsymbol{v}}^{\perp}\|^{2}/2$
$\leq H_{N}({\boldsymbol{\sigma}})-C\lambda^{2}N/3.$
@@
$\|{\boldsymbol{v}}^{i}_{\parallel}-{\boldsymbol{v}}^{j}_{\parallel}\|$ $=\|{\boldsymbol{v}}^{i}-{\boldsymbol{v}}^{j}\|{+}O(\|{\boldsymbol{v}}^{i}-{\boldsymbol{v}}^{i}_{\parallel}\|+\|{\boldsymbol{v}}^{j}-{\boldsymbol{v}}^{j}_{\parallel}\|)$
$=\|{\boldsymbol{\sigma}}^{i}-{\boldsymbol{\sigma}}^{j}\|{+}O(\|{\boldsymbol{v}}^{i}-{\boldsymbol{v}}^{i}_{\parallel}\|+\|{\boldsymbol{v}}^{j}-{\boldsymbol{v}}^{j}_{\parallel}\|)$
$=\Big{(}\sqrt{2-2(1-\varepsilon)^{2}}{+}O(\delta\varepsilon^{-1/2}+\lambda)\Big{)}\sqrt{N}.$
@@
The main part of the proof is the lower bound on ${\boldsymbol{\lambda}}_{K}(\nabla^{2}_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}_{\beta}))$. Suppose for sake of contradiction that (3.5) holds for some $K,C>0$, and define the event
$S_{{\rm eigen}}=\Big{\{}{\boldsymbol{\lambda}}_{K}(\nabla^{2}_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}_{\beta}))\leq-C\Big{\}}.$
Choose small $\varepsilon$ depending on $(K,C)$ such that $1-\varepsilon\in{\rm supp}(\zeta)$. Let $\eta\ll\delta$ be small depending on $(\xi,K,C,\varepsilon,\beta)$, and define for i.i.d. Gibbs samples $({\boldsymbol{\sigma}}_{\beta}^{1},\dots,{\boldsymbol{\sigma}}_{\beta}^{K+3})$ the event
@@
In light of Lemma 3.7 and the assumption $S_{{\rm bounded}}$, we find that $S_{{\rm eigen}}$ indeed cannot hold. Finally having shown (3.7), it follows that ${\mathbb{P}}[S_{{\rm eigen}}]\leq\frac{\varepsilon}{2}+e^{-cN}\leq\varepsilon$ which concludes the proof. ∎
@@ Acknowledgements
Thanks to Brice Huang for helpful discussions, and for comments on a previous version of the paper. We also thank Andrea Montanari, Pierfrancesco Urbani, and Lenka Zdeborová for enlightening conversations, as well as the anonymous referee for numerous helpful suggestions.